OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. Solve the system you wrote for this using any technique.

OpenStudy (saifoo.khan):

The perimeter of a triangle is just equal to the sum of the measures of the sides. Since the sides have measures x, x, and y, we know the perimeter is x+x+y, which is the same as 2x+y. Since the perimeter is equal to 46, our first equation is 2x+y=46. Since the measure y is 4 centimeters longer than the measure x, we know y is 4 more than x, so the second equation is y=x+4. So, the two equations are 2x+y=46 and y=x+4.

OpenStudy (saifoo.khan):

SAME?

OpenStudy (anonymous):

man you are a fast typer!

OpenStudy (anonymous):

faster than a speeding super khan

OpenStudy (saifoo.khan):

lol, i already wrote this in her last question.

OpenStudy (anonymous):

no. it's not! u need 2 solve the system for this and he copy-pasted!

OpenStudy (saifoo.khan):

i just copied from there.

OpenStudy (anonymous):

why do you need a system?

OpenStudy (saifoo.khan):

LOL. SUPER TYPER!

OpenStudy (anonymous):

oh because that is what it says. seems like a system is entirely unnecessary

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

really. you get \[x+x+x+4=46\] \[3x+4=46\] \[3x=42\] \[x=\frac{42}{3}=14\]

OpenStudy (anonymous):

Solve the system you wrote for this using any technique. that's what's different in this one!

OpenStudy (anonymous):

i just wrote it out yes? sides are x, x, and x+4 add together get 3x+4, which you know is 46. solve the linear equation and that is it. not much of a system

OpenStudy (saifoo.khan):

x = 14, y = 18